OpenStudy (anonymous):

Use the commutative property of addition to write an expression equivalent to the following: 9x+y

OpenStudy (zpupster):

commutative property of addition a+b = b+a 9x+y = y+9x

OpenStudy (anonymous):

@perl

OpenStudy (perl):

the only equivalent expression is y + 9x

OpenStudy (perl):

is there more information given